OpenStudy (anonymous):

the concentration of a certain drug in the bloodstream t hours after the drug is administered is given by c(t)=35te^(-bt) ng/ml, where b is some positive constant. suppose that the drug reaches maximum concentration exactly seven hours after being administered.Find the constant b.

OpenStudy (phi):

is this calculus ?

OpenStudy (anonymous):

Yes

OpenStudy (phi):

did you find the derivative of c(t) with respect to "t" ?

OpenStudy (phi):

is this calculus ?

OpenStudy (phi):

did you find the derivative of c(t) with respect to "t" ?

OpenStudy (anonymous):

-35e^(-bt) bt + 35e^(-bt)

OpenStudy (phi):

the standard technique to finding a min/max is to set the derivative to zero and in this case , solve for b

OpenStudy (phi):

and use drug reaches maximum concentration exactly seven hours set t= 7, so you get a numerical value for b

OpenStudy (phi):

the standard technique to finding a min/max is to set the derivative to zero and in this case , solve for b

OpenStudy (anonymous):

I did that and had something like -35e^(-7b) 7b + 35e^(-7b) which factors to (35e^(-7b)) (-7b+1)

OpenStudy (phi):

yes, and see that equal to 0

OpenStudy (anonymous):

-7b+1 =0 gives me b=0.1428 but i couldn't solve 35e^(-7b) = 0

OpenStudy (phi):

the standard technique to finding a min/max is to set the derivative to zero and in this case , solve for b

OpenStudy (phi):

yes, b= 1/7 the other factor can be ignored (in theory, b= infinity makes it approach zero, but we can't use that)

OpenStudy (phi):

Here is a graph of your function, with b= 1/7. As you can see, it peaks at x=7 (I used x instead t to plot it)

OpenStudy (anonymous):

ok thanks

OpenStudy (phi):

yw